A building in a city has a rectangular base. The length of the base measures
60 ft less than three times the width. The perimeter of this base is 960 ft. What are the dimensions of the​ base?
what is the width of base in feet? what is the length of base in feet?

Answer by addingup(3677)

L = 3W-60
2L + 2W = 960
Substitute for L:
2(3W-60)+2W = 960
Distribute the 2 (Distributive Properties of Multiplication)
6W-120+2W = 960
8W = 1080
W = 135
and the Length: (135*3)-60 = 345
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2(135) + 2(345) = 960
270 + 690 = 960 Correct